A small branch is wedged under a 200 kg rock and rests on a smaller object. The small object that creates a pivot point is called the fulcrum. The smaller object is 2.0 m from the large rock and the branch is 12.0 m long.
If the mass of the branch is negligible, what force must be exerted on the free end to just barely lift the rock?
1
Expert's answer
2020-01-20T05:18:37-0500


The above picture is the diagram illustrating the the above problem

we are given:

mass of the rock=200 kg

distance between the fulcrum and the rock= 2 m

distance between the fulcrum and the other end = (10-2)m= 8 m

we have to make the moment of all the forces must be equal to zero about the fulcrum

"\implies" moment of the big rock - moment of the applied force

"\implies" 200*10*2 - F*10 = 0

"\implies" F = 400 N